OpenStudy (anonymous):

write an equation of the line that passes through (1,6) and is parallel to the line 2y - 4x = 8

OpenStudy (anonymous):

Forming 2y-4x=8 to y=mx+b form, we get y=2x+4. y=mx+b Slopes are same for parallel lines, so: y=2x+b Plug in (1,6) to find your y-int. 6=2+b b=4 so your final equation for a parallel line that passes through (1,6) is: y=2x+4 and making that into ax+by form yields: y-2x=4